From njc@roe.ac.uk Fri Aug 12 17:56:59 2016 Date: Fri, 12 Aug 2016 16:56:48 +0000 From: Nicholas Cross To: Clive Davenhall , Bob Mann , Andy Lawrence , "Read, Mike" , Eckhard Sutorius , Mark S. Holliman , Nigel C. Hambly , Rob Blake , Ross Collins , Stelios Voutsinas , Dave Morris Cc: "dante@astro.puc.cl" , James Dunlop , "j.p.emerson@qmul.ac.uk" , ktn keithnoddle.org , "Lorenzo.Rimoldini@unige.ch" , "mcioni@aip.de" , "matt.jarvis@astro.ox.ac.uk" , "mike@ast.cam.ac.uk" , "nigel.metcalfe@durham.ac.uk" , "norman@astro.gla.ac.uk" , "pwl@star.herts.ac.uk" , "rgm@ast.cam.ac.uk" , "Simon.Dye@nottingham.ac.uk" , "s.j.warren@ic.ac.uk" , "tom.shanks@durham.ac.uk" Subject: WFAU VDFS Science Archives progress meeting minutes, 12 August 2016 WFAU VDFS Science Archives progress meeting minutes, 12 Aug 2016 Present: RPB, RSC, NJC, ETWS Apologies: ETWS, MAR, STV, DMR, KTN, AL, RGM, ACD, MSH DoNM: 10:00 am Friday 26th August 2016 in the Villas Meeting Room ACD in the chair (room is booked). Actions discharged this time: ------------------------------------------------ ACTION: RPB (was RPB, NJC) to determine if the necessary OGLE neighbour tables had been set-up and to ensure that they appear in the next VMC release. Discharged. Actions carried forward from 15/7/2016 meeting: ---------------------------------------------- ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan We currently have space so it was decided to start transferring DR12/DR13. Perhaps this action is now obsolete ACTION: ETWS to transfer SDSS DR12/13 Continues. ETWS had had contacted Ani Thakar, who had advised that SDSS DR13 would be released at the end of July (and that SDSS DR12 had recently been removed from the SDSS servers in preparation). RGM was happy to wait for DR13 but did not want further significant delays as one of his students planned to work with the data. ACTION: RPB, NJC to fix the errors encountered while trying to make non-survey releases U/14B/UA15, U/14B/UA18 and U/14B/UA19. Continues. RPB has rerun. U/14B/UA15 and U/14B/UA18 have completed. NJC to investigate errors still arising in U/14B/UA19. ACTION: ACD to liaise with DMR and STV to establish a set of deadlines for their work. Continues. ACTION: MAR to look at H-ATLAS documentation. Continues. NJC to discuss proposed VVVX meeting with RGM and to keep RSC and ACD apprised of developments pertaining to MOONS. NJC still to meet up with RGM .Continues Specific points and new actions: -------------------------------- Project management: * Nothing to report. WFCAM, VISTA and VST updates: * Nothing to report. Networking: * Nothing to report; the networks were operating nominally. WSA/VSA/OSA/GES Operations: * ETWS reported the following progress: - The VPHAS team release is reseaming at the moment (band merging complete), which will finish approx. end of October. * RPB reported the following progress: - 14B WFCAM non-surveys have been rerun, so moving onto new 15B - processing of the VMC P96 consortium release is at the variability processing. The release should be ready in ~10 days. - VVV is progressing and the TilePawPrints table is currently being updated and should take another month. * NJC reported the following progress: - The VVV ppErrorBit flags have been reset and the bug has been fixed. - VIKING ESO release files have been fixed. Hardware and Systems: * MSH noted that: - the copy of data off shepsesre is complete, so now I can begin working on rebuilding the disks. This is data from /disk39 and /disk40. - Ramses18 has arrived and is in a box in C2 awaiting installation and setup. Review of Software Release Deadlines: * Progress towards the software deadlines was reviewed briefly. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/PrepareSoft Other Software: * NJC decide that forced photometry from tiles was not working well, probably due to the poorer astrometric solution for tiles, as well as imperfections in grouting. He is reworking the pipeline to use pawprints and then averaging within the tiles. * ETWS noted that he is working on a script to find and delete focus files left in the transfer directories. These files should be deleted from the disks and the WSA DB. Review of Survey Release Deadlines: * Progress towards forthcoming survey deadlines was reviewed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Survey Data Release: * None Non-survey Data Release: * RPB reported that the following WSA 14B non-survey projects had been released: U/14B/UA15 and U/14B/UA18 Releases had been attempted but failed for a further 1 non-survey projects. The reason for these failures needs to be identified and rectified. ACTION: RPB, NJC. RPB should examine the log files to identify the problem and NJC or RPB should then fix it. Astrogrid deployment, VO & Data Analysis services: * Nothing to report. Miscellaneous: * NJC has registered for ADASS 2016 from October 16-20. NJC 12/8/16